Two identical particles,each of mass m and charge q,are seperated by a distance of 50cm in a medium of dielectric constant 81.If the magnitude of the electrostatic force is equal to the gravitational force between them,the specific charge of each particle in C/kg is

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Compare electrostatic force with gravitational force=mg

From here you can calculate the specific charge...

G. Navneeth
13 years ago

answer is  (81G/K)1/2 =Q/m=7.721x10-10 

since they r under equilibrium forces must be balance on equating those two forces will get the above answer..........from equating both the forces it is found that it is independent of distance ...which means they r under equilibrium at any position ,when the specific charges is as given above.